# Smart Contracts in the Real World: A Statistical Exploration of External Data Dependencies

### v3

@@ 4.2.6 Complementary
Data trustworthiness issues encompass both the inherent insecurity of the data itself and the possibility of data tampering, both of which can lead to the contamination of the data source. The execution of smart contracts relies on external data sources. However, an insecure data source may pose safety hazards to the smart contracts that depend on such data(Mao et al., 2024).
If the Oracle network fetches data from only one trading platform API, it cannot provide effective protection during that platform’s downtime, flash crashes, or price manipulation. This situation leads to the risk of smart contracts executing based on incorrect price data, potentially resulting in the loss of user funds. Data auditing is a primary method to ensure the security of data sources, but auditing external data by a third-party auditor (TPA) authorized by public auditing still presents security concerns, as TPA is often considered semi-honest. To address this issue, Kuan et al. proposed a decentralized auditing scheme (Dredas) by replacing the TPA with a designed smart contract. In this scheme, anyone can obtain audit results from Ethereum without worrying about a semi-honest TPA. Kuan’s article provides a solution to the external data trustworthiness issue by shifting data auditing from off-chain to on-chain. By designing smart contracts with data auditing capabilities, external data is audited on-chain before being utilized.
Oracles serve as a bridge for blockchains to access external data and multiple DeFi projects within the Ethereum network have encountered oracle issues, with some being attacked due to their reliance on a single oracle. Kevin et al. (Tjiam et al., 2021) have identified a manipulation process where attackers exploit unsecured flash loans to manipulate market prices provided by oracles, thereby illicitly profiting within smart contracts.
@@ 5 RELATED WORK
With the development of blockchain technology, the security of smart contracts has increasingly become a focal point. Several insightful studies have been proposed to investigate security issues within blockchain and smart contracts. Rafique et al. (Rafique and Qadir, 2024) provided a detailed introduction to the concept of Oracle services and elucidated the process through which smart contracts utilize the Oracle mechanism to obtain off-chain data. Dong et al. (Dong et al., 2023) point out that Oracle services always risk providing damaged, malicious, or inaccurate data. They introduce the Distributed Autonomous Oracle Network (DAON), its consensus protocol, and non-interactive schemes for reputation maintenance and payments. Vinayasree et al. (Vinayasree and Reddy, 2025) propose solving query performance issues and limited query semantics by linking to other databases. They achieve version control functionality by independently designing version control semantics. Luo et alċitep034. researched and describe widely used blockchain Oracle services, providing detailed insights into its potential roles, technical architecture, and design patterns.

### v4

@@ 4.2.6 Complementary
Data trustworthiness issues [39] encompass both the inherent insecurity of the data itself and the possibility of data tampering, both of which can lead to the contamination of the data source. The execution of smart contracts relies on external data sources. However, an insecure data source may pose safety hazards to the smart contracts that depend on such data [31].
If the Oracle network fetches data from only one trading platform API, it cannot provide effective protection during that platform’s downtime, flash crashes, or price manipulation. This situation leads to the risk of smart contracts executing based on incorrect price data, potentially resulting in the loss of user funds. Data auditing is a primary method to ensure the security of data sources, but auditing external data by a third-party auditor (TPA) [60, 33] authorized by public auditing still presents security concerns, as TPA is often considered semi-honest. To address this issue, Kuan et al. proposed a decentralized auditing scheme (Dredas) by replacing the TPA with a designed smart contract. In this scheme, anyone can obtain audit results from Ethereum without worrying about a semi-honest TPA. Kuan’s article provides a solution to the external data trustworthiness issue by shifting data auditing from off-chain to on-chain. By designing smart contracts with data auditing capabilities, external data is audited on-chain before being utilized.
Oracles serve as a bridge for blockchains to access external data and multiple DeFi projects within the Ethereum network have encountered oracle issues, with some being attacked due to their reliance on a single oracle. Kevin et al. [47] have identified a manipulation process where attackers exploit unsecured flash loans to manipulate market prices provided by oracles, thereby illicitly profiting within smart contracts.
@@ 5 RELATED WORK
With the development of blockchain technology, the security of smart contracts has increasingly become a focal point. Several insightful studies have been proposed to investigate security issues within blockchain and smart contracts. Rafique et al. [43] provided a detailed introduction to the concept of Oracle services and elucidated the process through which smart contracts utilize the Oracle mechanism to obtain off-chain data. Dong et al. [10] point out that Oracle services always risk providing damaged, malicious, or inaccurate data. They introduce the Distributed Autonomous Oracle Network (DAON), its consensus protocol, and non-interactive schemes for reputation maintenance and payments. Vinayasree et al. [50] propose solving query performance issues and limited query semantics by linking to other databases. They achieve version control functionality by independently designing version control semantics. Luo et al. [30]. researched and describe widely used blockchain Oracle services, providing detailed insights into its potential roles, technical architecture, and design patterns.